### v1

@@ Abstract
A subgraph $H$ of a graph $G$ is conformal if $G-V(H)$ has a perfect matching. This notion is intrinsically related to the aforementioned ear decomposition theorem — which implies that each matching covered graph (apart from $K_{2}$ and even cycles) contains a conformal bisubdivision of $\theta$, or a conformal bisubdivision of $K_{4}$, possibly both. (Here, $\theta$ refers to the graph with two vertices joined by three edges.) This immediately leads to two problems: characterize $\theta$-free (likewise, $K_{4}$-free) matching covered graphs. A characterization of planar $K_{4}$-free matching covered graphs was obtained by Kothari and Murty [J. Graph Theory, 82 (1), 2016]; the nonplanar case is open.
@@ 1 Introduction and summary
A graph is matchable if it has a perfect matching. A connected nontrivial graph is matching covered if each of its edges participates in some perfect matching. For instance, Schönberger (1927) proved that every $2$-connected cubic graph is matching covered; Figure 1 shows all such graphs of order at most six. This may also be deduced using Tutte’s $1$-factor Theorem stated below — where $\sf{c_{odd}}(H)$ denotes the number of odd components of a graph $H$.
@@
Figure 1: all $2$-connected cubic graphs of order six or less
Theorem 1.1.
@@
For distinct vertices $u$ and $v$ of a matchable graph $G$, the graph $G-u-v$ is not matchable if and only if there exists a barrier containing both $u$ and $v$. ∎
@@
A matchable graph is matching covered if and only if each of its barriers is stable; furthermore, if $B$ is a barrier of a matching covered graph $G$, then each component of $G-B$ is odd. ∎
@@
There is extensive literature on matching covered graphs; see [6]. Two important aspects of this theory are the ear decomposition theory and the tight cut decomposition theory, each of which plays an important role in our work; in the next section, we discuss the former that may be viewed as a refinement of Whitney’s ear decomposition theory for $2$-connected graphs.
@@ 1.1 Ear Decomposition Theory
Let $H$ be a subgraph of a graph $G$. We say that $H$ is a conformal subgraph if $G-V(H)$ is matchable. The following is easily proved by considering the symmetric difference of two perfect matchings.
@@
Clearly, one must exploit the symmetries of the Petersen graph $\mathbb{P}$. Since $\mathbb{P}$ has girth five and is non-hamiltonian, its only even cycles are $6$-cycles and $8$-cycles. If $C$ is any $8$-cycle, then $\mathbb{P}-V(C)$ is $K_{2}$; whereas if $C$ is any $6$-cycle, then $\mathbb{P}-V(C)$ is $K_{1,3}$. Thus, the conformal cycles of $\mathbb{P}$ are precisely its $8$-cycles; consequently, every conformal bisubdivision of $\theta$ (if one exists) must include an $8$-cycle as a subgraph. Figure 3 shows an $8$-cycle of $\mathbb{P}$, say $C$, in orange and red. The reader may observe that if $Q$ is any single ear of $C$ in $\mathbb{P}$, then $C+Q$ contains an odd cycle. This proves that $\mathbb{P}$ is indeed $\theta$-free.
Now, let $u$ and $v$ denote distinct vertices of $\mathbb{P}$. If $u$ and $v$ are adjacent, by Corollary 1.11, $\mathbb{P}+uv$ is $\theta$-based. Now suppose that $u$ and $v$ are nonadjacent; by exploiting the symmetries of $\mathbb{P}$, it suffices to consider one such pair. Observe that, in $\mathbb{P}$, there are two internally-disjoint odd $uv$-paths whose union is a conformal $8$-cycle $C$; thus $C+uv$ is a conformal bisubdivision of $\theta$ in $\mathbb{P}+uv$. This proves the second statement. ∎
@@ 1.2.1 Tight Cuts
We now proceed to state a characterization of bricks due to Edmonds, Lovász and Pulleyblank [7], which also implies (as per our discussion in the next Section 1.2.2) that the bricks and braces of a matching covered graph may be computed in poly-time. A matchable graph $G$ is bicritical if $G-u-v$ has a perfect matching for each pair $u,v\in V(G)$. A barrier is trivial if it comprises at most one vertex; otherwise it is nontrivial. Proposition 1.2 implies the following characterization.
@@
Bricks are precisely the $3$-connected bicritical graphs of order four or more.
In light of the above result, most authors including [9] require bicritical graphs to have at least four vertices. Note that, as per our relaxed definition, $K_{2}$ (up to multiple edges) is the only bicritical graph that is bipartite, and of course, the only one that has precisely two vertices. It is easily proved that every bicritical cubic graph is in fact $3$-connected; consequently, by the above theorem, each such graph, distinct from $\theta$, is a cubic brick. We remark that a matching covered graph of order four or more is bicritical if and only if its number of braces is zero.
@@ 1.2.2 ELP Cuts
Recall Corollary 1.3, and let $B$ denote a barrier of a matching covered graph $G$. Observe that if $L$ is any (odd) component of $G-B$, then $\partial_{G}(L)$ is a tight cut; such a tight cut is called a barrier cut. For an illustration, see the graphs shown in Figure 6. Observe that the graph $H$, obtained from $G$ by shrinking each component of $G-B$ to a single vertex, is a bipartite matching covered graph. This fact, coupled with Lovász’s Unique Tight Cut Decomposition Theorem (1.15), proves the following.
$C$$B$
@@
Let $B$ denote a barrier of a matching covered graph $G$, and let $\mathcal{L}$ denote the set of components of $G-B$. Then:
@@
Let $B$ denote a barrier of a matching covered graph $G$, and let $J$ denote any bicritical cubic graph. Then, $G$ is $J$-free if and only if, for each component $L$ of $G-B$, the graph $G/\overline{V(L)}$ is $J$-free.
@@
Figure 7: $2$-separation
Let $S:=\{u,v\}$ denote a $2$-vertex-cut, of a matching covered graph $G$, that is not a barrier; we refer to $S$ as a $2$-separation. It follows from Tutte’s Theorem (1.1) that $\sf{c_{odd}}(G-S)=0$. The reader may observe that, for any partition of $G-S$ into two disjoint subgraphs $G_{1}$ and $G_{2}$, the cuts $\partial_{G}(V(G_{1})\cup\{u\})$ and $\partial_{G}(V(G_{2})\cup\{u\})$ are nontrivial tight cuts; such tight cuts are called $2$-separation cuts. See the graphs shown in Figure 7 for illustrations.
Let us recall our discussion pertaining to the $\theta$-free graph $T_{6}$, shown in Figure 5, whose both $C$-contractions are $\theta$-based; this seems counter-intuitive in light of Theorem 1.18 that is applicable to all bicritical cubic graphs except $\theta$. To work around this issue, we use the following concept from Bondy and Murty [1]. For a $2$-vertex-cut $S:=\{u,v\}$ of a graph $G$, the graph $L^{\prime}:=G[V(L)\cup S]+uv$, where $L$ is a component of $G-S$, is called a marked $S$-component of $G$; the newly added edge joining $u$ and $v$ is called the marker edge of $L^{\prime}$. As shown in Figure 5, for $T_{6}$, both marked $S$-components are $K_{4}$ (without multiple edges) and thus $\theta$-free. We now make some other related observations.
Let $S:=\{u,v\}$ be a $2$-separation of a matching covered graph $G$, let $L$ denote a component of $G-S$, let $L^{\prime}$ denote the corresponding marked $S$-component, and let $L^{*}$ denote the tight cut contraction $G/(\overline{V(L)\cup\{u\}}\rightarrow v)$; see Figure 5. We invite the reader to compare $L^{\prime}$ and $L^{*}$, and observe that they are isomorphic except for the multiplicities of the edges joining $u$ and $v$ — each of which is at least one. This, coupled with Proposition 1.5, proves the following.
@@
Let $S:=\{u,v\}$ denote a $2$-separation of a matching covered graph $G$, let $L$ be a component of $G-S$, and let $L^{\prime}$ be the corresponding marked $S$-component. Then: (i) $L^{\prime}$ is matching covered, (ii) $L$ is matchable, and (iii) there exists a conformal (odd) $uv$-path in $L$. ∎
In light of Proposition 1.14, if $J$ is any simple matching covered graph, then $L^{\prime}$ is $J$-free if and only if $L^{*}$ is $J$-free. However, since our primary interest, namely $\theta$, is not simple, we must resort to marked $S$-components (instead of tight cut contractions) in the case of $2$-separation cuts. Also, the above observations, coupled with Lovász’s Unique Tight Cut Decomposition Theorem (1.15), imply the following.
@@
Let $S$ denote a $2$-separation of a matching covered graph $G$, and let $\mathcal{L^{\prime}}$ denote the set of marked $S$-components of $G$. Then:
@@
Let $S$ denote a $2$-separation of a matching covered graph $G$. If $G-S$ has three or more components, then $G$ is $\theta$-based. ∎
In light of the above statement, the technical assumption in the next statement is quite harmless, and in fact necessary. Its proof appears in Section 2.2.
@@
[Recursing across a $2$-separation]
Let $S$ denote a $2$-separation of a matching covered graph $G$, and let $J$ denote any $3$-connected cubic graph. Assume that if $J=\theta$ then $G-S$ has precisely two components. Then, $G$ is $J$-free if and only if each marked $S$-component of $G$ is $J$-free.
@@
Every matching covered graph, that has a nontrivial tight cut, either has a nontrivial barrier cut or a $2$-separation cut, possibly both.
In light of the above, by an ELP cut, we mean a nontrivial tight cut that is either a barrier cut or a $2$-separation cut. It is worth noting that a matching covered graph may have a nontrivial tight cut that is not an ELP cut. For instance, the reader may verify that the tight cut $C$, shown in Figure 4, is not an ELP cut; however, the cut $D$ is an ELP cut. The reader may find it surprising that the original proof of the above theorem relied on linear programming duality. Szigeti [18] gave the first graph-theoretical proof. More recently, Carvalho, Lucchesi and Murty [5] gave yet another proof that relies heavily on the theory developed by them. As alluded to earlier, the tight cut decomposition of a matching covered graph is poly-time computable. This is precisely due to the Canonical Partition Theorem (1.4) and the ELP Theorem (1.26).
@@
    $G$ has a nontrivial barrier and for any such barrier, say $B$, the graph $G/\overline{V(L)}$ is $\theta$-free for each component $L$ of $G-B$, or
@@
    $G$ has a $2$-separation and for any such $2$-separation, say $S$, the graph $G-S$ has precisely two components, and both marked $S$-components of $G$ are $\theta$-free. ∎
Our above theorem implies a poly-time algorithm to decide whether a given matching covered graph $G$ is $\theta$-free; we describe this briefly. If $G$ is a brick or a brace, we simply use statement (i). Otherwise, we compute the canonical partition of $G$; see Theorem 1.4. If $G$ has a maximal barrier $B$ that is nontrivial, we recurse on the $|B|$ smaller graphs as defined in statement (ii). Otherwise, $G$ is bicritical; see Corollary 1.16. By the ELP Theorem (1.26), $G$ has a $2$-separation cut. We compute any $2$-vertex-cut $S$ of $G$; since $G$ is bicritical, $S$ is a $2$-separation. As per statement (iii), if $G-S$ has three or more components, we conclude that $G$ is $\theta$-based; otherwise, we recurse on the two marked $S$-components of $G$. We refer the reader to [6] for a discussion on the facts that the (a) canonical partition may be computed in poly-time and (b) checking whether a graph is a brick or a brace is poly-time. We leave the runtime analysis to the reader.
@@
Now suppose that statement (ii) of Theorem 1.29 holds, and let $B$ denote a nontrivial barrier. Observe that, by Corollary 1.16, $G$ is not bicritical; thus $G\notin\mathcal{T}$. For each $1\leq i\leq|B|$, let $L_{i}$ denote the components of $G-B$, and for each $i$, we let $L^{\prime}_{i}:=G/\overline{V(L_{i})}$, and let $n_{i},m_{i}$ and $b_{i}$ denote the order, size and the number of bricks, respectively, of $L^{\prime}_{i}$. Since each $L^{\prime}_{i}$ is a smaller $\theta$-free graph, by the induction hypothesis, $m_{i}\leq\frac{3n_{i}}{2}+b_{i}-1$ and $b_{i}\leq\frac{n_{i}}{2}-1$. Now, we use easy counting arguments and Proposition 1.20 to conclude that $G$ satisfies the claimed bounds with strict inequality:
@@
Finally, suppose that statement (iii) of Theorem 1.29 holds, and let $S$ denote a $2$-separation. We let $L^{\prime}_{1}$ and $L^{\prime}_{2}$ denote the marked $S$-components of $G$; for each $i\in\{1,2\}$, let $n_{i},m_{i}$ and $b_{i}$ denote the order, size and the number of bricks, respectively, of $L^{\prime}_{i}$. Since both $L^{\prime}_{i}$s are smaller $\theta$-free graphs, by the induction hypothesis, $m_{i}\leq\frac{3n_{i}}{2}+b_{i}-1$, and equality holds if and only if $L^{\prime}_{i}\in\mathcal{T}$; furthermore, $b_{i}\leq\frac{n_{i}}{2}-1$, and equality holds if and only if $L^{\prime}_{i}\in\mathcal{T}_{0}$. Since each $L^{\prime}_{i}$ has order four or more, by Corollary 1.11, each of them is simple; consequently, $S$ is a stable set in $G$. Note that $m=m_{1}+m_{2}-2$ and $n=n_{1}+n_{2}-2$. By easy counting arguments and Proposition 1.23, we conclude that the bound in statement (i) holds:
@@
We remark that the $K_{2}$-sum operation that appears in Corollary 1.30 is reminiscent of the $C_{4}$-sum (aka $4$-cycle sum) operation that was crucial in the solution of Problem 4 for bipartite graphs — in the works of Robertson, Seymour and Thomas [17], and of McCuaig [16].

### v2

@@ Abstract
A subgraph $H$ of a graph $G$ is conformal if GV(H)G-V(H) has a perfect matching. This notion is intrinsically related to the aforementioned ear decomposition theorem — which implies that each matching covered graph (apart from $K_{2}$ and even cycles) contains a conformal bisubdivision of $\theta$, or a conformal bisubdivision of $K_{4}$, possibly both. (Here, $\theta$ refers to the graph with two vertices joined by three edges.) This immediately leads to two problems: characterize $\theta$-free (likewise, $K_{4}$-free) matching covered graphs. A characterization of planar $K_{4}$-free matching covered graphs was obtained by Kothari and Murty [J. Graph Theory, 82 (1), 2016]; the nonplanar case is open.
@@ 1 Introduction and summary
A graph is matchable if it has a perfect matching. A connected nontrivial graph is matching covered if each of its edges participates in some perfect matching. For instance, Schönberger (1927) proved that every 22-connected cubic graph is matching covered; Figure 1 shows all such graphs of order at most six. This may also be deduced using Tutte’s 11-factor Theorem stated below — where $\sf{c_{odd}}(H)$ denotes the number of odd components of a graph $H$.
@@
Figure 1: all 22-connected cubic graphs of order six or less
Theorem $1.$
@@
For distinct vertices $u$ and $v$ of a matchable graph $G$, the graph GuvG-u-v is not matchable if and only if there exists a barrier containing both $u$ and $v$. ∎
@@
A matchable graph is matching covered if and only if each of its barriers is stable; furthermore, if $B$ is a barrier of a matching covered graph $G$, then each component of GBG-B is odd. ∎
@@
There is extensive literature on matching covered graphs; see [6]. Two important aspects of this theory are the ear decomposition theory and the tight cut decomposition theory, each of which plays an important role in our work; in the next section, we discuss the former that may be viewed as a refinement of Whitney’s ear decomposition theory for 22-connected graphs.
@@ 1.1 Ear Decomposition Theory
Let $H$ be a subgraph of a graph $G$. We say that $H$ is a conformal subgraph if GV(H)G-V(H) is matchable. The following is easily proved by considering the symmetric difference of two perfect matchings.
@@
Clearly, one must exploit the symmetries of the Petersen graph $\mathbb{P}$. Since $\mathbb{P}$ has girth five and is non-hamiltonian, its only even cycles are 66-cycles and 88-cycles. If $C$ is any 88-cycle, then $\mathbb{P}-V(C)$ is $K_{2}$; whereas if $C$ is any 66-cycle, then $\mathbb{P}-V(C)$ is $K_{1,3}$. Thus, the conformal cycles of $\mathbb{P}$ are precisely its 88-cycles; consequently, every conformal bisubdivision of $\theta$ (if one exists) must include an 88-cycle as a subgraph. Figure 3 shows an 88-cycle of $\mathbb{P}$, say $C$, in orange and red. The reader may observe that if $Q$ is any single ear of $C$ in $\mathbb{P}$, then $C+Q$ contains an odd cycle. This proves that $\mathbb{P}$ is indeed $\theta$-free.
Now, let $u$ and $v$ denote distinct vertices of $\mathbb{P}$. If $u$ and $v$ are adjacent, by Corollary 1.11, $\mathbb{P}+uv$ is $\theta$-based. Now suppose that $u$ and $v$ are nonadjacent; by exploiting the symmetries of $\mathbb{P}$, it suffices to consider one such pair. Observe that, in $\mathbb{P}$, there are two internally-disjoint odd uvuv-paths whose union is a conformal 88-cycle $C$; thus $C+uv$ is a conformal bisubdivision of $\theta$ in $\mathbb{P}+uv$. This proves the second statement. ∎
@@ 1.2.1 Tight Cuts
We now proceed to state a characterization of bricks due to Edmonds, Lovász and Pulleyblank [7], which also implies (as per our discussion in the next Section 1.2.2) that the bricks and braces of a matching covered graph may be computed in poly-time. A matchable graph $G$ is bicritical if GuvG-u-v has a perfect matching for each pair $u,v\in V(G)$. A barrier is trivial if it comprises at most one vertex; otherwise it is nontrivial. Proposition 1.2 implies the following characterization.
@@
Bricks are precisely the 33-connected bicritical graphs of order four or more.
In light of the above result, most authors including [9] require bicritical graphs to have at least four vertices. Note that, as per our relaxed definition, $K_{2}$ (up to multiple edges) is the only bicritical graph that is bipartite, and of course, the only one that has precisely two vertices. It is easily proved that every bicritical cubic graph is in fact 33-connected; consequently, by the above theorem, each such graph, distinct from $\theta$, is a cubic brick. We remark that a matching covered graph of order four or more is bicritical if and only if its number of braces is zero.
@@ 1.2.2 ELP Cuts
Recall Corollary 1.3, and let $B$ denote a barrier of a matching covered graph $G$. Observe that if $L$ is any (odd) component of GBG-B, then $\partial_{G}(L)$ is a tight cut; such a tight cut is called a barrier cut. For an illustration, see the graphs shown in Figure 6. Observe that the graph $H$, obtained from $G$ by shrinking each component of GBG-B to a single vertex, is a bipartite matching covered graph. This fact, coupled with Lovász’s Unique Tight Cut Decomposition Theorem (1.15), proves the following.
CCBB
@@
Let $B$ denote a barrier of a matching covered graph $G$, and let $\mathcal{L}$ denote the set of components of GBG-B. Then:
@@
Let $B$ denote a barrier of a matching covered graph $G$, and let $J$ denote any bicritical cubic graph. Then, $G$ is $J$-free if and only if, for each component $L$ of GBG-B, the graph $G/\overline{V(L)}$ is $J$-free.
@@
Figure 7: 22-separation
Let $S:=\{u,v\}$ denote a 22-vertex-cut, of a matching covered graph $G$, that is not a barrier; we refer to $S$ as a 22-separation. It follows from Tutte’s Theorem (1.1) that $\sf{c_{odd}}(G-S)=0$. The reader may observe that, for any partition of GSG-S into two disjoint subgraphs $G_{1}$ and $G_{2}$, the cuts $\partial_{G}(V(G_{1})\cup\{u\})$ and $\partial_{G}(V(G_{2})\cup\{u\})$ are nontrivial tight cuts; such tight cuts are called 22-separation cuts. See the graphs shown in Figure 7 for illustrations.
Let us recall our discussion pertaining to the $\theta$-free graph $T_{6}$, shown in Figure 5, whose both $C$-contractions are $\theta$-based; this seems counter-intuitive in light of Theorem 1.18 that is applicable to all bicritical cubic graphs except $\theta$. To work around this issue, we use the following concept from Bondy and Murty [1]. For a 22-vertex-cut $S:=\{u,v\}$ of a graph $G$, the graph $L^{\prime}:=G[V(L)\cup S]+uv$, where $L$ is a component of GSG-S, is called a marked $S$-component of $G$; the newly added edge joining $u$ and $v$ is called the marker edge of $L^{\prime}$. As shown in Figure 5, for $T_{6}$, both marked $S$-components are $K_{4}$ (without multiple edges) and thus $\theta$-free. We now make some other related observations.
Let $S:=\{u,v\}$ be a 22-separation of a matching covered graph $G$, let $L$ denote a component of GSG-S, let $L^{\prime}$ denote the corresponding marked $S$-component, and let $L^{*}$ denote the tight cut contraction $G/(\overline{V(L)\cup\{u\}}\rightarrow v)$; see Figure 5. We invite the reader to compare $L^{\prime}$ and $L^{*}$, and observe that they are isomorphic except for the multiplicities of the edges joining $u$ and $v$ — each of which is at least one. This, coupled with Proposition 1.5, proves the following.
@@
Let $S:=\{u,v\}$ denote a 22-separation of a matching covered graph $G$, let $L$ be a component of GSG-S, and let $L^{\prime}$ be the corresponding marked $S$-component. Then: (i) $L^{\prime}$ is matching covered, (ii) $L$ is matchable, and (iii) there exists a conformal (odd) uvuv-path in $L$. ∎
In light of Proposition 1.14, if $J$ is any simple matching covered graph, then $L^{\prime}$ is $J$-free if and only if $L^{*}$ is $J$-free. However, since our primary interest, namely $\theta$, is not simple, we must resort to marked $S$-components (instead of tight cut contractions) in the case of 22-separation cuts. Also, the above observations, coupled with Lovász’s Unique Tight Cut Decomposition Theorem (1.15), imply the following.
@@
Let $S$ denote a 22-separation of a matching covered graph $G$, and let $\mathcal{L^{\prime}}$ denote the set of marked $S$-components of $G$. Then:
@@
Let $S$ denote a 22-separation of a matching covered graph $G$. If GSG-S has three or more components, then $G$ is $\theta$-based. ∎
In light of the above statement, the technical assumption in the next statement is quite harmless, and in fact necessary. Its proof appears in Section $2.$
@@
[Recursing across a 22-separation]
Let $S$ denote a 22-separation of a matching covered graph $G$, and let $J$ denote any 33-connected cubic graph. Assume that if $J=\theta$ then GSG-S has precisely two components. Then, $G$ is $J$-free if and only if each marked $S$-component of $G$ is $J$-free.
@@
Every matching covered graph, that has a nontrivial tight cut, either has a nontrivial barrier cut or a 22-separation cut, possibly both.
In light of the above, by an ELP cut, we mean a nontrivial tight cut that is either a barrier cut or a 22-separation cut. It is worth noting that a matching covered graph may have a nontrivial tight cut that is not an ELP cut. For instance, the reader may verify that the tight cut $C$, shown in Figure 4, is not an ELP cut; however, the cut $D$ is an ELP cut. The reader may find it surprising that the original proof of the above theorem relied on linear programming duality. Szigeti [18] gave the first graph-theoretical proof. More recently, Carvalho, Lucchesi and Murty [5] gave yet another proof that relies heavily on the theory developed by them. As alluded to earlier, the tight cut decomposition of a matching covered graph is poly-time computable. This is precisely due to the Canonical Partition Theorem (1.4) and the ELP Theorem (1.26).
@@
    $G$ has a nontrivial barrier and for any such barrier, say $B$, the graph $G/\overline{V(L)}$ is $\theta$-free for each component $L$ of GBG-B, or
@@
    $G$ has a 22-separation and for any such 22-separation, say $S$, the graph GSG-S has precisely two components, and both marked $S$-components of $G$ are $\theta$-free. ∎
Our above theorem implies a poly-time algorithm to decide whether a given matching covered graph $G$ is $\theta$-free; we describe this briefly. If $G$ is a brick or a brace, we simply use statement (i). Otherwise, we compute the canonical partition of $G$; see Theorem 1.4. If $G$ has a maximal barrier $B$ that is nontrivial, we recurse on the $|B|$ smaller graphs as defined in statement (ii). Otherwise, $G$ is bicritical; see Corollary 1.16. By the ELP Theorem (1.26), $G$ has a 22-separation cut. We compute any 22-vertex-cut $S$ of $G$; since $G$ is bicritical, $S$ is a 22-separation. As per statement (iii), if GSG-S has three or more components, we conclude that $G$ is $\theta$-based; otherwise, we recurse on the two marked $S$-components of $G$. We refer the reader to [6] for a discussion on the facts that the (a) canonical partition may be computed in poly-time and (b) checking whether a graph is a brick or a brace is poly-time. We leave the runtime analysis to the reader.
@@
Now suppose that statement (ii) of Theorem 1.29 holds, and let $B$ denote a nontrivial barrier. Observe that, by Corollary 1.16, $G$ is not bicritical; thus $G\notin\mathcal{T}$. For each $1\leq i\leq|B|$, let $L_{i}$ denote the components of GBG-B, and for each $i$, we let $L^{\prime}_{i}:=G/\overline{V(L_{i})}$, and let $n_{i},m_{i}$ and $b_{i}$ denote the order, size and the number of bricks, respectively, of $L^{\prime}_{i}$. Since each $L^{\prime}_{i}$ is a smaller $\theta$-free graph, by the induction hypothesis, $m_{i}\leq\frac{3n_{i}}{2}+b_{i}-1$ and $b_{i}\leq\frac{n_{i}}{2}-1$. Now, we use easy counting arguments and Proposition 1.20 to conclude that $G$ satisfies the claimed bounds with strict inequality:
@@
Finally, suppose that statement (iii) of Theorem 1.29 holds, and let $S$ denote a 22-separation. We let $L^{\prime}_{1}$ and $L^{\prime}_{2}$ denote the marked $S$-components of $G$; for each $i\in\{1,2\}$, let $n_{i},m_{i}$ and $b_{i}$ denote the order, size and the number of bricks, respectively, of $L^{\prime}_{i}$. Since both $L^{\prime}_{i}$s are smaller $\theta$-free graphs, by the induction hypothesis, $m_{i}\leq\frac{3n_{i}}{2}+b_{i}-1$, and equality holds if and only if $L^{\prime}_{i}\in\mathcal{T}$; furthermore, $b_{i}\leq\frac{n_{i}}{2}-1$, and equality holds if and only if $L^{\prime}_{i}\in\mathcal{T}_{0}$. Since each $L^{\prime}_{i}$ has order four or more, by Corollary 1.11, each of them is simple; consequently, $S$ is a stable set in $G$. Note that $m=m_{1}+m_{2}-2$ and $n=n_{1}+n_{2}-2$. By easy counting arguments and Proposition 1.23, we conclude that the bound in statement (i) holds:
@@
We remark that the $K_{2}$-sum operation that appears in Corollary 1.30 is reminiscent of the $C_{4}$-sum (aka 44-cycle sum) operation that was crucial in the solution of Problem 4 for bipartite graphs — in the works of Robertson, Seymour and Thomas [17], and of McCuaig [16].